OpenStudy (anonymous):

Convert the radian measure to degree measure. Use the value of π found on a calculator and round answers to two decimal places.

OpenStudy (anonymous):

9pie/4

OpenStudy (anonymous):

You should have pi on your calculator as a symbol, if not just use pi=3.1415. Now you have the value, \[\frac{9 \pi}{4}\]and we know that the factor to convert from radian to degree is\[\frac{180}{\pi}\]So when you multiply the two you get,\[\frac{9 \pi}{4} \times \frac{180}{\pi}\]Now multiply the fractions and you get the result. Hint: The pi is simplified.

OpenStudy (anonymous):

do you do cross product?

OpenStudy (anonymous):

Well, the best way to solve this is like:\[\frac{9*180}{4}\]